xx CarpetReduce

CarpetReduce

Erik Schnetter <schnetter@uni-tuebingen.de>

April 29, 2003

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11debugxiweight 10Scope: privatex01 BOOLEAN

11Description: Allow debugging iweight grid function by keeping it allocated

11                      10  x01  Default: no

11

11minxmaxxtimexinterpolation 10Scope: privatex01  BOOLEAN

11Description: Interpolate in time for min/max reductions

11                      10  x01  Default: yes

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: Produce screen output while running

11                      10  x01  Default: no

11

7 Interfaces

General Implements: reduce

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 iweight 10 iweight x01 compact110
11                 10  x01 description11Integer weight mask
11 10 x x01 description11using 2D̂ bits
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11 10 x01 tags11prolongation=”none” InterpNumTimelevels=1 checkpoint=”no”
11                 10  x01 timelevels111
11                 10  x01variable type11INT
11 weight 10 weight x01 compact110
11                 10  x01 description11Weight function
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11 10 x01 tags11prolongation=”none” InterpNumTimelevels=1 checkpoint=”no”
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 one 10 one x01 compact110
11                 10  x01 description11Constant one
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11 10 x01 tags11prolongation=”none” InterpNumTimelevels=1 checkpoint=”no”
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 excisedxcells 10 excisedxcells x01 compact110
11                 10  x01 description11Excised (ignored) volume in the simulation domain
11 10 x x01 description11in terms of coarse grid cells
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11 10 x01 tags11checkpoint=”no”
11                 10  x01 timelevels111
11                 10  x01variable type11REAL

11

Adds header: bits.h to CarpetReducexbits.h

Uses header: nompi.h defs.hh dh.hh dist.hh vect.hh carpet.hh carpet.h typecase.hh typeprops.hh loopcontrol.h

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/CarpetReduce. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always: 11x
11weight  11x
11excisedxcells11x

11x  11x

11

Scheduled Functions

CCTKxSTARTUP x carpetreducestartup x startup routine x

11x11Language:11c
11 x 11Type: 11function

11

CCTKxBASEGRID x maskbasexsetupmask x set up the weight function x

11x11After:11spatialcoordinates
11 x 11x  11sphericalsurfacexsetup
11 x 11Type:11group

11

MaskBasexSetupMaskAll x maskbasexsetmask x set the weight function x

11x11After: 11setupimask
11 x 11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Reads: 11carpetreduce::iweight(everywhere)
11 x 11Type: 11function
11 x 11Writes: 11carpetreduce::one(everywhere)
11 x 11x  11carpetreduce::weight(everywhere)

11

MaskBasexSetupMaskAll x setupmask x set up the real weight function (schedule other routines in here)

11x11After:11maskbasexsetmask
11 x 11Type:11group

11

MaskBasexSetupMaskAll x maskbasextestmask x test the weight function x

11x11After: 11setupmask
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Reads: 11carpetreduce::excisedxcells(everywhere)
11 x 11Type: 11function
11 x 11Writes: 11carpetreduce::excisedxcells(everywhere)

11

SetupIMaskInternal x coordbasexsetupmask x set up the outer boundaries of the weight function x

11x11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Type: 11function
11 x 11Writes: 11carpetreduce::iweight(everywhere)

11

SetupIMaskInternal x carpetmasksetup x set up the weight function for the restriction regions x

11x11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-singlemap
11 x 11Reads: 11carpetreduce::iweight(everywhere)
11 x 11Type: 11function
11 x 11Writes: 11carpetreduce::iweight(everywhere)

11

CCTKxPOSTREGRIDINITIAL x maskbasexsetupmask x set up the weight function x

11x11After:11spatialcoordinates
11 x 11Type:11group

11

CCTKxPOSTREGRID x maskbasexsetupmask x set up the weight function

11x11After:11spatialcoordinates
11 x 11Type:11group

11

CCTKxPOSTxRECOVERxVARIABLES x maskbasexsetupmask x set up the weight function x

11x11Type:11group

11

MaskBasexSetupMask x maskbasexsetupmaskall x set up the weight function x

11x11Type:11group

11

MaskBasexSetupMaskAll x maskbasexallocatemask x allocate the weight function x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function
11 x 11Writes: 11carpetreduce::excisedxcells(everywhere)

11

MaskBasexSetupMaskAll x maskbasexinitmask x initialise the weight function x

11x11After: 11maskbasexallocatemask
11 x 11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Type: 11function
11 x 11Writes: 11carpetreduce::iweight(everywhere)

11

MaskBasexSetupMaskAll x setupimaskinternal x set up the integer weight function (schedule other routines in here) x

11x11After:11maskbasexinitmask
11 x 11Type:11group

11

MaskBasexSetupMaskAll x setupimask x set up the integer weight function (schedule other routines in here) x

11x11After:11setupimaskinternal
11 x 11Type:11group